Lucas Company has provided the following information: Cash flow from operating activities, $361,400 Net income, $308,800 Interest expense, $32,100 Interest cash payments, $21,400 Income tax payments, $254,000 Income tax expense, $247,400 What was Lucas' times interest earned ratio?

29.76

18.33

27.30

18.20
Answer #1

Answer:18.33 Times

Times Interest Earned measures the ability of the company to pay off its debt payments .it is also called the Interest coverage ratio.

Times Interest Earned = [ Net income + Interest expense + Income tax expense ] / Interest expense

Times Interest Earned = [ 308800 + 32100 + 247400 ] / 32100 = 588300 / 32100 = 18.33 Times